OpenStudy (anonymous):

The graph of which function will have a maximum and a y-intercept of 4?

OpenStudy (anonymous):

A) f(x) = 4x2 + 6x – 1 B) f(x) = –4x2 + 8x + 5 C) f(x) = –x2 + 2x + 4 D) f(x) = x2 + 4x – 4

OpenStudy (anonymous):

@Elsa213

Elsa213 (elsa213):

What do you think the answer is?

OpenStudy (anonymous):

i was thinkin about B

undeadknight26 (undeadknight26):

B has a maximum of 9.

OpenStudy (anonymous):

so am i right? 0_0

Elsa213 (elsa213):

I think so cx

undeadknight26 (undeadknight26):

I dunno ._.

Elsa213 (elsa213):

/.\

undeadknight26 (undeadknight26):

Im assuming its 4x^2 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i forgot to put that

undeadknight26 (undeadknight26):

And a maximum of 4 ._.

OpenStudy (anonymous):

A) f(x) = 4x^2 + 6x – 1 B) f(x) = –4x^2 + 8x + 5 C) f(x) = –x^2 + 2x + 4 D) f(x) = x^2 + 4x – 4

undeadknight26 (undeadknight26):

Why would you say B? To me it's none of these ._.

OpenStudy (anonymous):

let me check real quick hold on

OpenStudy (anonymous):

i pick C it's correct

OpenStudy (anonymous):

i don'tknow how to example

undeadknight26 (undeadknight26):

C is correct?

undeadknight26 (undeadknight26):

I would say it would be also...Although technically the maximum is of it is actually 5 it's the only one that actually crosses 0,4.